OpenStudy (anonymous):

Theo wants to use a cookie recipe that makes 36 cookies, but he wants to reduce the number if cookies to 24. If the recipe specifies 2 cups of sugar, how much sugar should he use?

OpenStudy (anonymous):

1.3333 cups 24/36 = X/2 X = 2/3 * 2 = 1.3333

OpenStudy (esshotwired):

1 1/3 cups because you will have to divide 36/24 which equals 1.5. Then to find out how much sugar you need for 24 cookies, you have to divide 2/1.5. This equals 1.333 but it also equals 1/3 cups